RC meeting and Run Plan, Thursday February 9th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Report


  • Waiting most of the last day for machine optics to get us the right beam size.
  • Good beam width finally achieved an production is on to 45 uA on He3.
  • BBCal shows some blocks without data

Current Status

  • Running 45 uA on He3

Run Plan

  • Continue running 45 uA on He3

Data Collection and Polarization Status

  • 24.0C of uncorrected charge for this kinematic (26.4C of corrected charge)
  • Last NMR 50%

Other Topics/Status Updates

  • BBCal Currently being looked into.
  • Anything else?

Shift Schedule

  • Shift Schedule Page
  • Only 9 shift spots open between owl 17th Feb and owl March 20th. All open slots are TO - please encourage someone to sign up - especially someone who only has 6 shifts!

Requested Accesses to the Hall

  • Time of next access opportunity currently not clear
  • Jack would like to collect some cables
  • GEM team need like a 3-4hour access for the SBS GEMs DAQ troubleshooting